OpenStudy (anonymous):

The area of a square garden is 128 m². How long is the diagonal?

OpenStudy (austinl):

 \(A_{square}=l^2\) \(128=l^2\) \(d^2=l^2+l^2\) Can you solve from there?

OpenStudy (anonymous):

hello my fellow people im barack ombama and i have an important message for you

OpenStudy (anonymous):

nope i dont no how

OpenStudy (austinl):

\(l^2=128\) \(d^2=l^2+l^2~=~2\times l^2~\Leftrightarrow~2\times128=\) \(d^2=256\) \(d=\sqrt{256}=~?\)

OpenStudy (anonymous):

so confusing

OpenStudy (anonymous):

:/

OpenStudy (anonymous):